The force exerted by the wind on the sails of a sailboat is Fsail = 300 N north. The water exerts a force of Fkeel = 240 N east. If the boat (including its crew) has a mass of 260 kg, what are the magnitude and direction of its acceleration?The force exerted by the wind on the sails of a sa
Answer #1

Fr = sqrt(Fx^2 + Fy^2)

Fr = sqrt(240^2 + 300^2)

Fr = 384 N

direction = tan^-1(Fy/Fx) = tan^-1(300/240) = 51.3 degree north of east

resultant acceleration of the boat is = Fr/m = 384/260 = 1.48 m/s^2

direction = 51.3 degree north of east